20 EN/M2 Frd Fices al A, B, and c 2 2
Answer #1

Give load is uniformly distributed load.

Assuming unit width in the direction perpendicular to the plane of the paper we get load intensity (w) = 20 * 1 = 20 kN/m per unit width.

Now force at a point located at depth Z is given by,

Fz = w * Z

Therefore

Force at A = 20 * 2 = 40 kN per unit width

Force at B = 20 * (2 + 2) = 80 kN per unit width

Force at C = 20 * (2 + 2 + 1) = 100 kN per unit width.